OpenStudy (anonymous):

help!! please!! ill give you a medal. What is the equation in point-slope form of the line passing through (2, 2) and (4, 8)?

OpenStudy (here_to_help15):

@phi

OpenStudy (anonymous):

Point slope form is y - y1 = m(x-x1). We can use any set of those points but we need the slope first, and we can find that by y2-y1/x2-x1. The slope is 6/2 or 3. Now we can use point slope form. Let's use the first coordinates. y - 2 = 3(x-2) Just distribute the right side first... y -2 = 3x-6 and add 2 to isolate y. y = 3x-4.